# CLaMP 3: Universal Music Information Retrieval Across Unaligned Modalities and Unseen Languages

### v2

@@ 3.2 Metadata Curation
RAG: Using Qwen2.5-72B Yang et al. (2024), we generated annotations from the retrieved documents and basic metadata (titles and artist names). We refined the metadata when discrepancies were found. The annotations covered the fields in Table 1 under Annotations, with an additional Boolean field indicating if the source material had sufficient information for generating meaningful annotations.
@@ 3.4 Benchmark Dataset
Appendix B provides t-SNE visualizations of CLaMP 3 embeddings on WikiMT-X, showing modality, language, and semantic distributions in the shared representation space. In addition, Table 11 presents the genre classification results of CLaMP 3 and baseline models across different musical modalities and text annotations.
@@ 4 Experiments
This section evaluates CLaMP 3 on retrieval tasks, comparing it to state-of-the-art baselines. We present results for the two best-performing CLaMP 3 variants—one for symbolic music and one for audio. A full retrieval comparison of all variants can be found in Appendix C, and classification results are available in Appendix D.
@@ 4.2 English Text-to-Music Retrieval
For symbolic music retrieval, we compared CLaMP 3 with CLaMP 2 Wu et al. (2024) and CLaMP Wu et al. (2023a) on WikiMT (using ABC notation) and MidiCaps Melechovsky et al. (2024) (using MIDI). For audio retrieval, we evaluated CLaMP 3 against state-of-the-art models CLAP Wu et al. (2023b) and TTMR++ Doh et al. (2024) on the Song Describer Dataset (SDD) Manco et al. (2023) and MusicCaps-Remake (MC-R) Agostinelli et al. (2023), which addresses data leakage by using full-length audio and rewritten captions (see Appendix E) from AudioSet’s evaluation set Gemmeke et al. (2017). In addition, we tested all models on WikiMT-X to evaluate their performance across varying semantic perspectives.
@@ 4.3 Multilingual Text-to-Music Retrieval
CLaMP 3’s two variants differ in their language exposure. CLaMP 3 initializes its text and symbolic music encoders from CLaMP 2, which was pre-trained on symbolic-text alignment across all XLM-R-supported languages, giving it prior exposure to all languages in Table 3. In contrast, CLaMP 3 has never aligned music data with the languages marked with asterisks, demonstrating true cross-modal generalization in its performance.
@@ 6 Limitations
In addition, we would like to express our gratitude for the use of icons from flaticon333https://www.flaticon.com in Fig. 1 and Fig. 2. We also thank Yusong Wu from the University of Montreal for helping us understand details of CLAP.
@@ Appendix A Prompt and Examples
Appendix B t-SNE Visualizations on WikiMT-X
@@
Fig. 8(c) shows four semantic categories—Background, Analysis, Description, and Scene—showing how CLaMP 3 handles different content types. Background, Analysis, and Description often converge, reflecting the overlap in explanatory texts as they cover related musical concepts. In contrast, Scene forms distinct clusters, likely because it focuses on visual depictions, leading to more consistent semantic patterns tied to specific imagery rather than music.
@@
Appendix C Performance of CLaMP 3 Variants
@@
Appendix D Music Classification
@@
D.1 Symbolic Music Classification
@@
D.2 Audio Music Classification
@@
Table 9 shows the strengths of contrastive and self-supervised models in audio classification. CLaMP 3 variants excel in high-level tasks, like genre classification (GTZAN) and tagging (MTT), where capturing abstract musical meaning is crucial. MERT, however, performs better in low-level tasks such as key detection (GS) and pitch classification (Nsynth), where fine spectral detail is more important. Contrastive models generally struggle with short-duration audio (e.g., 4-second clips in Nsynth) because their focus on aligning longer segments with text limits their ability to capture fine acoustic details. These results suggest contrastive learning is better for semantic tasks, while self-supervised models are more effective for detailed acoustic analysis, particularly for short clips.
@@ D.3 Classification on WikiMT-X
Appendix E Data Leakage of MusicCaps

### v3

@@ 3.2 Metadata Curation
RAG: Using Qwen2.5-72B Yang et al. (2024), we generated annotations from the retrieved documents and basic metadata (titles and artist names). The annotations covered the fields in Table 1 under Annotations, with an additional Boolean field indicating if the source material had sufficient information for generating meaningful annotations.
@@ 3.4 Benchmark Dataset
Appendix B presents detailed objective and human evaluations of WikiMT-X annotation quality. In addition, Appendix C provides t-SNE visualizations of CLaMP 3 embeddings on WikiMT-X, showing modality, language, and semantic distributions in the shared representation space.
@@ 4 Experiments
This section evaluates CLaMP 3 on retrieval tasks, comparing it to state-of-the-art baselines. We present results for the two best-performing CLaMP 3 variants—one for symbolic music and one for audio. A full retrieval comparison of all variants can be found in Appendix D, and classification results are available in Appendix E.
@@ 4.2 English Text-to-Music Retrieval
For symbolic music retrieval, we compared CLaMP 3 with CLaMP 2 Wu et al. (2024) and CLaMP Wu et al. (2023a) on WikiMT (using ABC notation) and MidiCaps Melechovsky et al. (2024) (using MIDI). For audio retrieval, we evaluated CLaMP 3 against state-of-the-art models CLAP Wu et al. (2023b) and TTMR++ Doh et al. (2024) on the Song Describer Dataset (SDD) Manco et al. (2023) and MusicCaps-Remake (MC-R) Agostinelli et al. (2023), which addresses data leakage by using full-length audio and rewritten captions (see Appendix F) from AudioSet’s evaluation set Gemmeke et al. (2017). In addition, we tested all models on WikiMT-X to evaluate their performance across varying semantic perspectives.
@@ 4.3 Multilingual Text-to-Music Retrieval
CLaMP 3’s two variants differ in their language exposure. CLaMP 3 initializes its text and symbolic music encoders from CLaMP 2, which was pre-trained on symbolic-text alignment across all XLM-R-supported languages, giving it prior exposure to all languages in Table 3. In contrast, CLaMP 3 has never aligned music data with the languages marked with asterisks, demonstrating true cross-lingual generalization in its performance.
@@ 6 Limitations
In addition, we thank Flaticon333https://www.flaticon.com for icons used in Fig. 1 and Fig. 2, Yusong Wu (University of Montreal) for helping us understand CLAP in detail, and Monan Zhou (Central Conservatory of Music) for assisting with WikiMT-X data processing.
@@ Appendix A Prompt and Examples
| Dataset | Annotation | CLaMP | CLaMP 2 | CLaMP 3 | CLaMP 3 | Alignment | Aesthetics |
| WikiMT | Caption | 0.1900 | 0.1244 | 0.2028 | 0.2184 | 4.83 | 3.42 |
| MidiCaps | Caption | 0.1133 | 0.0255 | 0.1401 | 0.1583 | 3.92 | 2.83 |
| WikiMT-X | Background | 0.2429 | 0.1343 | 0.2239 | 0.2264 | 4.67 | 3.50 |
| Analysis | 0.1336 | 0.1097 | 0.2261 | 0.2461 | 3.75 | ||
| Description | 0.0794 | 0.0451 | 0.2359 | 0.2752 | 3.42 | ||
| Scene | 0.0779 | 0.1410 | 0.2240 | 0.2874 | 3.67 | ||
| Dataset | Annotation | CLAP | TTMR++ | CLaMP 3 | CLaMP 3 | Alignment | Aesthetics |
| SDD | Caption | 0.3446 | 0.3340 | 0.1129 | 0.1683 | 2.25 | 2.50 |
| MusicCaps | Caption | 0.2518 | 0.3957 | 0.1077 | 0.1500 | 4.08 | 2.58 |
| WikiMT-X | Background | 0.3734 | 0.4477 | 0.0092 | 0.1186 | 4.50 | 3.58 |
| Analysis | 0.2594 | 0.3813 | 0.0024 | 0.1298 | 3.92 | ||
| Description | 0.2000 | 0.3340 | 0.0385 | 0.1738 | 3.33 | ||
| Scene | 0.1848 | 0.2590 | 0.0525 | 0.1996 | 3.83 |
Appendix B Evaluation of Annotation Quality Across Benchmarks
To evaluate the semantic quality of music-text pairings, we conduct an evaluation combining (i) automatic similarity scores from models and (ii) human ratings on music-text alignment and musical aesthetics. The results are summarized in Table 5.
We include both LLM-generated and human-annotated datasets in this evaluation. Specifically, all symbolic datasets—WikiMT, MidiCaps, and WikiMT-X—contain text annotations generated by large language models. In contrast, the audio datasets differ in annotation quality: MusicCaps captions are written by trained musicians, while SDD was annotated by non-expert crowd workers.
For the automatic evaluation, we report the average cosine similarity between text and music embeddings generated by multiple retrieval models, including CLAP, TTMR++, CLaMP, CLaMP 2, and two CLaMP 3 variants. These similarity scores indicate how well the text and music are aligned in the shared embedding space.
WikiMT-X consistently achieves higher similarity scores across most annotation types compared to previous datasets. For example, its Background annotations reach cosine similarities of 0.2429 (CLaMP), 0.3734 (CLAP), and 0.4477 (TTMR++), outperforming scores seen in other datasets. Similar improvements are observed in the Analysis and Description categories, with the highest overall score of 0.2874 achieved by CLaMP 3 on the Scene annotation type.
To complement the automatic metrics, we conducted a human evaluation. Four conservatory-trained musicians (each with over 10 years of formal experience) rated 144 music-text examples based on two criteria: semantic alignment (how well the text matches the music) and musical aesthetics (the perceived quality of the music), using a 1-5 scale.
Results show that WikiMT-X performs on par with or better than expert-annotated datasets. In the symbolic domain, its Background annotations receive a high alignment score of 4.67—close to WikiMT’s 4.83 and higher than MidiCaps’ 3.92.
In the audio domain, WikiMT-X again demonstrates strong performance. Its alignment scores reach up to 4.50, and aesthetics up to 3.58—substantially outperforming SDD (2.25 / 2.50) and MusicCaps (4.08 / 2.58), which include more amateur or crowd-sourced material. The consistently higher aesthetic ratings for WikiMT and WikiMT-X likely stem from their inclusion of well-known Western popular music from the 20th century, in contrast to the less polished recordings found in other benchmarks.
Overall, the results support WikiMT-X as a high-quality benchmark for multimodal text-to-music retrieval. Its annotations show robust semantic alignment and musical relevance, confirmed by both model metrics and expert ratings. With careful filtering and validation, LLM-generated annotations can match or even exceed expert quality across both symbolic and audio datasets.
Appendix C t-SNE Visualizations on WikiMT-X
@@ Appendix B Evaluation of Annotation Quality Across Benchmarks
Fig. 8(c) shows four semantic categories—Background, Analysis, Description, and Scene—presenting how CLaMP 3 handles different content types. Background, Analysis, and Description often converge, reflecting the overlap in explanatory texts as they cover related musical concepts. In contrast, Scene forms distinct clusters, likely because it focuses on visual depictions, leading to more consistent semantic patterns tied to specific imagery rather than music.
@@
Appendix D Performance of CLaMP 3 Variants
@@
Appendix E Music Classification
@@
E.1 Symbolic Music Classification
@@
E.2 Audio Music Classification
@@
Table 10 shows the strengths of contrastive and self-supervised models on the MARBLE benchmark. CLaMP 3 variants excel in high-level tasks, like genre classification (GTZAN) and tagging (MTT), where capturing abstract musical meaning is crucial. MERT, however, performs better in low-level tasks such as key detection (GS) and pitch classification (Nsynth), where fine spectral detail is more important. Contrastive models generally struggle with short-duration audio (e.g., 4-second clips in Nsynth) because their focus on aligning longer segments with text limits their ability to capture acoustic details. These results suggest contrastive learning is better for semantic tasks, while self-supervised models are more effective for low-level acoustic analysis.
@@ E.3 Classification on WikiMT-X
Appendix F Data Leakage of MusicCaps